The Square of a positive number is 12 more than 4 times the number. Find the number.
Answer by ankor@dixie-net.com(22740)

The Square of a positive number is 12 more than 4 times the number. Find the number.
:
x^2 = 4x + 12
x^2 - 4x - 12 = 0, our old friend the quadratic equation
Factors to
(x-6)(x+2) = 0
Positive solution
x = 6
:
:
check: 6^2 = 4(6) + 12